OpenStudy (anonymous):

Find equation of the tangent line to the curve x^(2)-2(sqrtxy)-y^(3)==11 at the point (4,1)

OpenStudy (turingtest):

\[x^2-2\sqrt{xy}-y^3=11\]?

OpenStudy (turingtest):

^is this your problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know how to implicitly differentiate?

OpenStudy (anonymous):

If someone could double check my algebra..I calculated the derivative as \[(dy/dx)=(2x-(\sqrt(y)/\sqrt(x)))/(3y^2-(\sqrt(x)/\sqrt(y)))\]